# Adaptive Uncertainty Quantification for Generative AI

### v1

@@ Abstract
This supplementary material contains the proof of all results as well as auxiliary lemmata in Section A. Section B includes a description of a modified algorithm that achieves tighter theoretical guarantees at the cost of making computation scale linearly in the number of test points. In Section C, we provide implementation details for all experiments and simulations present in the article, including data sources and preprocessing information, hyperparameter choices, and LLM prompt information. We also include additional visualizations referred to in the text in Section C.3.
@@ 1 Introduction
Conformal prediction [37] is arguably one of the most widely used tools for predictive uncertainty quantification. We develop a conformal prediction method in the context of contemporary applications involving black-box predictive systems (such as generative models including ChatGPT [4], BERT [8], LLaMA [34], and Watson Assistant [9]), which have been pre-trained on massive datasets that are obscured from the user.
@@
In the absence of the training dataset, split-conformal prediction (isolating fitting from ranking) is a viable alternative [18, 19] to full conformal prediction as it avoids model refitting on augmented training data. Split-conformal prediction ranks conformity scores on the calibration dataset and guarantees finite-sample marginal coverage without distributional assumptions or any knowledge of the actual mechanism inside the pre-trained model. While standard conformal prediction bands are marginal over the covariates, and thereby overly conservative, a conditional coverage guarantee is not achievable in general [16, 36, 2]. This has motivated the development of various refinements that achieve some relaxed form of conditional coverage [36, 13, 35] such as group-conformal prediction [17, 2].
@@
Locally adaptive conformal prediction bands have been sought that shrink and stretch depending on the difficulty of the prediction problem at each local point [19, 27, 6]. In particular, Lei et al. [19] locally enhance conformal prediction by rescaling the conformity scores with a function fitted on the absolute residuals of the training data. This approach is not feasible without access to training data and, moreover, can lead to underestimation of the prediction error when the black-box model has overfitted the training data [27]. Alternatively, Romano et al. [27] conformalize quantile estimates so that the tightness of the prediction bands depends on the tightness (conformity) of the quantile black-box model in use. Rossellini et al. [29] further build on this by incorporating the uncertainty in the quantile estimation into the prediction intervals. Chernozhukov et al. [6] generalize this conformalized quantile regression as a fully quantile-rank-based method by using a probability integral transform with a distributional regression estimator (a model of the conditional CDF). These approaches require an auxiliary function (residuals, quantiles, or CDFs), which is obtained using the training dataset. In this work, we concern ourselves with the modern scenario where the users are provided a basic black-box predictive model (without quantiles) and do not have access to the training data.
@@
Another way of tightening prediction intervals is through conditioning on a finite set of groups. Lei and Wasserman [16] apply conformal prediction to each local set for kernel density estimation and study its asymptotic conditional validity. The locality is achieved by an equally spaced partition of the data space, where the data is bisected: one part is used for tuning the spacing size, and the other for model fitting and conformal prediction. Our approach allows for non-equispaced partitions which are capable of (local) adaptation [26, 5]. When a good partition is already known, one can consider existing works such as [13, 2]. Barber et al. [2] consider a countable set of potentially overlapping groups with sufficiently large probability mass. The prediction band for a data point is determined by the largest prediction band obtained by applying conformal prediction to each group that contains the data point. Gibbs et al. [13] reformulate conditional coverage as coverage over covariate shifts. When the probability of the test data is tilted by an indicator function of a subset in the covariate space, this guarantee becomes equivalent to a set conditional coverage guarantee.
@@
Our work is built upon these previous efforts in the sense that we promote local adaptivity through the lens of group-conditioned conformal prediction. Unlike most existing works [13, 2], however, we do not assume that the groups are known in advance. We enhance group-conformal inference by using data-adaptive grouping driven by a partitioning model fitted to the conformity scores of the calibration set. Enabling self-grouping in conditional conformal prediction with continuous predictors is a step forward towards improving local adaptivity in a wider range of settings. Our self-grouping is cast as a tree-based supervised learning where each group is represented by a bin (box) of a box-shaped partition of the covariate space. We then apply the standard split-conformal prediction inside each newfound bin. Equivalently, this translates as constructing the prediction intervals by adding an entire step function as opposed to only a constant (as would be the case in split-conformal). Motivated by the fact that (dyadic) trees can adapt to varying levels of local smoothness [26], our procedure (called Conformal Tree) deploys dyadic trees for locally adaptive prediction band construction. Since the training set is not available in our scenario, we cannot re-calibrate conformity scores by fitting a (tree or other) model to the conformity scores on the training data (as was done in [19]). A naive strategy would be to bisect the calibration set and to use one half for conformity score construction and the second for calibration. Our same-set approach allows us to avoid further sample splitting and use the entire calibration dataset. We account for, and protect against, any systematic differences between calibration and training residuals that might occur due to covariate shifts or model fitting strategies (e.g. intentional overfitting by interpolation). Our predictive interval nearly maintains the desired group-conditional coverage level (See, Lemma 5 and Theorem 3.1), paying a negligible price in exchange for the ability to re-use the calibration data. Second, compared to a full-conformal-style prediction band that applies the partition mechanism on the union of the calibration set and the test data points (and thereby would have to be re-fitted for all values of the training conformity score), our robust tree approach allows fitting the partition model only once. Our computational approach is different from a related approach by Gibbs et al. [13] who perform conditional quantile estimation over certain function classes, avoiding refitting on augmented data by exploiting monotonicity properties and dual formulation of quantile regression.
@@
Exchangeability among the calibration and test data points is the key assumption that makes conformal prediction free from distributional assumptions, ensuring uniformity of the ranks of conformity scores. However, conditioning on groups learned only from the calibration set, excluding the test data, may break down exchangeability. To obviate the need to refit the partition model for any possible realization of testing data, we propose a modification of the classical CART algorithm. First, we allow only for splits at dyadic rationals (not at observed values) that are not subject to shift after adding more data. Second, we propose a robust range-based split criterion and a stopping rule based on a node size. We show that adding a new observation changes the estimated partition only with a small probability controlled by a chosen tuning parameter. This property allows us to conclude a group-conditional coverage lower bound at a level which only slightly (indistinguishably in practice) deviates from the chosen coverage level.
@@ 2 Conformal Tree
The core idea behind our self-grouping conformal prediction method is to learn the group information by applying a tree model to the conformity scores of the calibration dataset (Section 2.1). Since this may break the exchangeability assumption between the test and calibration data points, we develop a tree algorithm that is robust to adding a new test observation (Section 2.2).
@@ 2.1 Calibration by Self-grouping
Note that the prediction interval construction relies on the model’s performance in the overall input space. For example, with the absolute residual score $S(x,y)=|y-\hat{\mu}(x)|$, the set $\widehat{C}_{n}(X_{n+1})$ is equivalent to $[\hat{\mu}(X_{n+1})-\hat{q},\hat{\mu}(X_{n+1})+\hat{q}]$, where a constant $\hat{q}=S^{*}(\mathcal{D}_{1:n})$ is added and subtracted over all the data domain $\mathcal{X}$.
To sharpen the prediction interval, we can consider adding and subtracting a constant that differs depending on the location of $X_{n+1}$. Assume that the data space $\mathcal{X}=\cup_{k=1}^{K}\mathscr{X}_{k}$ is partitioned into non-overlapping groups $\mathfrak{X}=\{\mathscr{X}_{1},\ldots,\mathscr{X}_{K}\}$. We allow for categorical covariates which naturally serve as a group indicator. If $X_{n+1}\in\mathscr{X}_{k}$, we can calibrate the conformal interval based on only data points inside $\mathscr{X}_{k}$ to obtain a group-conditional coverage guarantee. In other words, defining
@@
Our method, called Conformal Tree, is described inside Algorithm 1. This procedure has an additional partition step compared to group-conformal inference [13, 2] because we do not assume that the grouping information is known a priori. We apply a regression tree on the conformity scores of the calibration dataset $\mathcal{D}_{1:n}$ using a potentially high-dimensional vector of both continuous and discrete covariates. Denote by $\widehat{\mathcal{T}}(\mathcal{D}_{1:n})\in\mathbb{T}_{K}$ the tree fitted on $\mathcal{D}_{1:n}$, where $\mathbb{T}_{K}$ is the class of binary dyadic trees with $K$ leaf nodes (defined formally later). The partition induced by the tree $\widehat{\mathcal{T}}(\mathcal{D}_{1:n})$ is denoted by $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n})$, where calibration is performed locally through (2.1) with $\mathfrak{X}=\widehat{\mathfrak{X}}(\mathcal{D}_{1:n})$. The fact that grouping information depends only on the calibration data, and not the test point, can potentially lead to a violation of group-conditional exchangeability for the conformity scores (see, Lemma 2). However, if we were to obtain a partition $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$ based on the union of calibration and testing data $\mathcal{D}_{1:n+1}$, the triples $\left(X_{1},Y_{1},S(X_{1},Y_{1})\right),...,\left(X_{n+1},Y_{n+1},S(X_{n+1},Y_{n+1})\right)$ would still be exchangeable when conditioned on $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$. Motivated by this observation, we propose a grouping mechanism based on a robust tree approach (Section 2.2), which guarantees that the partition does not change too often after adding one more observation, i.e. $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n})=\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$ with high probability. Such "unchangeability" property guarantees that the conditional coverage (that would be guaranteed if $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$ were used for splitting) is inherited with high probability when $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n})$ is used instead.
@@
1:Fitted model $\hat{\mu}$, calibration dataset $\mathcal{D}_{1:n}$, tree hyperparameters ($m,K,\eta=0.05$)
@@ 2.2 Robust Regression Trees
We focus on dyadic trees where splits along covariates may only be placed at specific predetermined locations (dyadic rationals for continuous predictors rescaled to a unit interval). In general, our implementation allows for discrete predictors where dummy variable are created and the split occurs at 1/2. Ordinal discrete covariates are rescaled and treated as continuous.
We define a binary tree $\mathcal{T}$ as a collection of hierarchically organized nodes $(l,k)$ such that $(l,k)\in\mathcal{T}\Rightarrow(j,\lceil k/2^{l-j}\rceil)\in\mathcal{T}$ for $j=0,...,l-1$. Here $l$ denotes the depth, $k$ denotes the horizontal position, and the root node is defined to be $(0,0)$. The left and right children of a node $(l,k)$ are given by $(l+1,2k)$ and $(l+1,2k+1)$, respectively. A node $(l,k)\in\mathcal{T}$ is called an external or leaf node if it has no children. We denote the set of external nodes of $\mathcal{T}$ by $\mathcal{T}_{\rm leaves}$ and the internal nodes $\mathcal{T}_{\rm int}=\mathcal{T}\setminus\mathcal{T}_{\rm leaves}$. The internal nodes are assigned a split direction while the leaf nodes are not. We denote these split directions by $\mathcal{H}(\mathcal{T})=\{h_{lk}\in\{1,\dots,d\}:(l,k)\in\mathcal{T}_{int}\}$. We consider only dyadic trees, for which the tree nodes correspond to axis-aligned dyadic splits. We define $\mathbb{T}_{K}$ as a space of all dyadic binary trees that split at most $K-1$ times in one of the $d$ directions yielding up to $K$ leaves. The partition $\mathfrak{X}(\mathcal{T})$, induced by both $\mathcal{T}$ and $\mathcal{H}(\mathcal{T})$, consists of regions associated with the leaf nodes $\mathcal{T}_{\rm leaves}$. In particular, each leaf node $(l,k)\in\mathcal{T}_{\rm leaves}$ corresponds to a rectangular region of the input space $\mathcal{X}$, which we denote as $\mathcal{X}_{(l,k)}(\mathcal{T},\mathcal{H}(\mathcal{T}))$. Given data $X_{1},\ldots,X_{n}\in\mathcal{X}$ and scores $S_{1},\ldots,S_{n}\in\mathbb{R}$, we denote the subset of data points that fall inside the $(l,k)^{th}$ region as $X_{(l,k)}(\mathcal{T},\mathcal{H}(\mathcal{T}))=\{X_{i}:X_{i}\in\mathcal{X}_{(l,k)}(\mathcal{T},\mathcal{H}(\mathcal{T}))\}$ and, similarly, $S_{(l,k)}(\mathcal{T},\mathcal{H}(\mathcal{T}))=\{S_{i}:X_{i}\in\mathcal{X}_{(l,k)}(\mathcal{T},\mathcal{H}(\mathcal{T}))\}$.
Our robust tree-fitting criterion is based on a splitting rule that targets the response range (as opposed to deviance). At every split opportunity, we split at the terminal node in the direction that yields the largest range decrease. Below, we formally define the range reduction of a node $(l,k)$ if it were split dyadically along the dimension $j$ denoted by $I_{l,k}^{j}(\mathcal{T})$.
@@
For a tree $\mathcal{T}\in\mathbb{T}_{K}$ with split directions $\mathcal{H}(\mathcal{T})$, the range of a node $(l,k)\in\mathcal{T}_{\rm leaves}$ is given by $R_{lk}(\mathcal{T})=\max S_{(l,k)}(\mathcal{T},\mathcal{H}(\mathcal{T}))-\min S_{(l,k)}(\mathcal{T},\mathcal{H}(\mathcal{T}))\,.$ When a node $(l,k)$ splits along the $j^{\textrm{th}}$ direction, we define the range of the left child as $R^{j,\rm left}_{lk}(\mathcal{T})=\max S_{(l+1,2k)}(\widetilde{\mathcal{T}},\mathcal{H}^{j})-\min S_{(l+1,2k)}(\widetilde{\mathcal{T}},\mathcal{H}^{j})\,,$ where $\widetilde{\mathcal{T}}=\mathcal{T}\cup\{(l+1,2k),(l+1,2k+1)\}$ and $\mathcal{H}^{j}=\mathcal{H}(\mathcal{T})\cup\{h_{lk}=j\}.$ The range of the right child $R^{j,\rm right}_{lk}(\mathcal{T})$ is defined similarly as $S_{(l+1,2k+1)}(\widetilde{\mathcal{T}},\mathcal{H}^{j})$. Then, the range reduction obtained by splitting a node $(l,k)$ in dimension $j$ is given by
$I_{lk}^{j}(\mathcal{T})=R_{lk}(\mathcal{T})-\left[R^{j,\rm left}_{lk}(\mathcal{T})+R^{j,\rm right}_{lk}(\mathcal{T})\right]\big{/}2\,.$ (2.2)
As is typical with CART implementations [25], we regularize the tree and define a stopping criterion by the following three hyperparameters: the minimal number of samples per leaf $m$, maximum number of leaves $K$, and minimum range reduction rate $\eta=0.05$. A node is called a candidate node if there exists a direction along which a split would not cause the tree to violate any of these criteria.
@@
We use the notation from Definition 2.1. For a tree $\mathcal{T}\in\mathbb{T}_{K}$ such that $|\mathcal{T}_{\rm leaves}|<K$, the set of eligible split directions for a node $(l,k)\in\mathcal{T}_{\rm leaves}$ is defined as
@@
We continue to sequentially add splits to the tree by maximizing the range reduction among candidate nodes until either we reach the maximum number of leaves $K$, or there are no candidates remaining, at which time we stop the tree fitting procedure.
@@
1:Dataset $\mathcal{D}=\{(X_{i},S_{i}):i\in[n]\}$, minimum samples per leaf $m$, max leaves $K$
@@
3:while $|\mathcal{T}_{\mathrm{leaves}}|<K$  do
@@
Our robust tree algorithm resembles greedy CART [3]. Similarly, we require a minimum improvement (range reduction rate) and a minimum number of observations per leaf. Rather than using cross-validation to prune back, we rely on early stopping once we reach $K$ leaf nodes or until splitting would yield less than $m$ observations per leaf node, or an insufficient improvement. The tuning parameter $K$ is pre-determined, unlike Bayesian implementations of CART [7] which are both locally and globally adaptive.

### v2

@@ Abstract
This supplementary material contains the proof of all results as well as auxiliary lemmata in Section A. Section B includes a description of a modified algorithm that achieves tighter theoretical guarantees at the cost of making computation scale linearly in the number of test points. Section C conducts sensitivity analyses for the main hyperparameters $K_{\rm max}$ and $m$. In Section D, we provide implementation details for the experiments and simulations present in the article, including data sources and preprocessing information, hyperparameter choices, and LLM prompt information. We also include additional experiments and visualizations referred to in the text in Section E.
@@ 1 Introduction
Artificial Intelligence (AI) continues to permeate swiftly through all facets of society, including high-stakes decision domains such as healthcare [12, 64, 2], justice [11, 50, 46], and/or finance [56, 35]. The practical utility of AI in decision-making depends critically on the ability of the underlying foundational model to effectively communicate uncertainty.
Large language models (LLMs) may be able to convey uncertainty through repeated prompting [63, 49], self-consistency checks [60], adversarially perturbed inputs [65], or by simply asking LLM to verbalize its own uncertainty [32, 52, 63]. While these heuristic approaches (see [21] for more references) seek uncertainty quantification through internal reproducibility checks, we focus on external evaluations through a labeled calibration dataset. In particular, this work develops a predictive uncertainty quantification method in the context of contemporary applications involving black-box predictive systems (such as generative models including ChatGPT [5], BERT [13], LLaMA [54], and Watson Assistant [14]) based on conformal prediction.
@@
Conformal prediction [58] is arguably one of the most widely used tools for quantification of predictive uncertainty. It has been used in the context of LLMs to provide statistical guarantees for the factuality of model responses [8, 36, 31]. In our work, we revisit the more traditional statistical frameworks of regression and classification in situations where a proprietary labeled dataset is available to calibrate predictions from a model trained on data that are unavailable to the user. In the absence of the training dataset, split-conformal prediction (isolating fitting from ranking) is a viable alternative to full conformal prediction [28, 29] as it avoids model refitting on augmented training data. Split-conformal prediction ranks conformity scores on the calibration dataset and guarantees finite-sample marginal coverage without distributional assumptions or any knowledge of the actual mechanism inside the pre-trained model. While standard conformal prediction bands are marginal over the covariates, and thereby overly conservative in many local regions, a conditional coverage guarantee is not achievable in general [26, 57, 3]. This has motivated the development of various refinements that achieve some relaxed form of conditional coverage [57, 18, 55] such as group-conformal prediction [27, 3].
@@
Locally adaptive conformal prediction bands have been sought that shrink and stretch depending on the difficulty of the prediction problem at each local point [29, 42, 9]. In particular, Lei et al. [29] locally enhance conformal prediction by rescaling the conformity scores with a function fitted on the absolute residuals of the training data. This approach is not feasible without access to training data and, moreover, can lead to underestimation of the prediction error when the black-box model has overfitted the training data [42]. Alternatively, Romano et al. [42] conformalize quantile estimates so that the tightness of the prediction bands depends on the tightness (conformity) of the quantile black-box model in use. Rossellini et al. [44] further build on this by incorporating the uncertainty in the quantile estimation into the prediction intervals. Chernozhukov et al. [9] generalize this conformalized quantile regression as a fully quantile-rank-based method by using a probability integral transform with a distributional regression estimator (a model of the conditional CDF). These approaches require an auxiliary function (residuals, quantiles, or CDFs), which is obtained using the training dataset. In this work, we concern ourselves with the modern scenario where the users are provided a basic black-box predictive model (without quantiles) and do not have access to the training data. Note that this setup aligns with the modern foundational models whose massive training datasets are obscured from the user.
@@
Another way of tightening prediction intervals is through conditioning on a finite set of groups. Lei and Wasserman [26] apply conformal prediction to each local set for kernel density estimation and study its asymptotic conditional validity. The locality is achieved by an equally spaced partition of the data space, where the data is bisected: one part is used for tuning the spacing size, and the other for model fitting and conformal prediction. Our approach allows for non-equispaced partitions which are capable of (local) adaptation [41, 7]. When a good partition is already known, one can consider existing methods such as those described by [18, 3]. Barber et al. [3] consider a countable set of potentially overlapping groups with sufficiently large probability mass. The prediction band for a data point is determined by the largest prediction band obtained by applying conformal prediction to each group that contains the data point. Gibbs et al. [18] reformulate conditional coverage as coverage over covariate shifts. When the probability of the test data is tilted by an indicator function of a subset in the covariate space, this guarantee becomes equivalent to a set conditional coverage guarantee.
@@
Our work is built upon these previous efforts in the sense that we promote local adaptivity through the lens of group-conditioned conformal prediction. In this work, we do not assume that the groups are known in advance. We enhance group-conformal inference by using data-adaptive grouping driven by a partitioning model fitted to the conformity scores of the calibration set. Enabling self-grouping in conditional conformal prediction with continuous predictors is a step forward towards improving local adaptivity in a wider range of settings. Our self-grouping is cast as a tree-based supervised learning where each group is represented by a bin (box) of a box-shaped partition of the covariate space. We then apply the standard split-conformal prediction inside each newfound bin. Equivalently, this translates as constructing the prediction intervals by adding an entire step function as opposed to only a constant (as would be the case in split-conformal). Motivated by the fact that (dyadic) trees can adapt to varying levels of local smoothness [41], our procedure (called Conformal Tree) deploys dyadic trees for locally adaptive prediction band construction. Since the training set is not available in our scenario, we cannot re-calibrate conformity scores by fitting a (tree or other) model to the conformity scores on the training data (as was done in [29]).
@@
Exchangeability among the calibration and test data points is the key assumption that ensures validity of conformal prediction sets without requiring distributional assumptions, by ensuring uniformity of the ranks of conformity scores. However, conditioning on groups learned only from the calibration set, excluding the test data, may no longer ensure such an exchangeability property. To obviate the need to refit the partition model for any possible realization of testing data, we propose a modification of the classical CART algorithm. First, we allow only for splits at dyadic rationals (not at observed values) that are not subject to shift after adding more data. Second, we propose a robust range-based split criterion and a stopping rule based on a node’s size. We show that adding a new observation changes the estimated partition only with a small probability controlled by a chosen tuning parameter. This property allows us to conclude a group-conditional coverage lower bound at a level which only slightly (indistinguishably in practice) deviates from the chosen coverage level. The distribution-free aspect distinguishes our work from [62, 24, 20], who developed local conformal prediction methods under suitable distributional assumptions. After the completion of this manuscript, several related independent approaches were brought to our attention [33, 23, 6]. The approach Martinez et al. [33] is most similar to our work in that it also fits a regression tree on the conformity scores. However, our same-set approach is unique as it uses the entire calibration dataset for both calibration and partitioning. To achieve grouping, all of these previous works require additional data on top of the calibration data. The driving force behind our same-set approach is our new partitioning tree method that exhibits a unique features that we call "unchangeability", i.e. robustness to adding a new observation. Our same-set approach is particularly appealing when the calibration dataset is not overwhelmingly large
We account for, and protect against, any systematic differences between calibration and training residuals that might occur due to covariate shifts or model fitting strategies (e.g. intentional overfitting by interpolation). Our predictive interval nearly maintains the desired group-conditional coverage level (see Lemma 5 and Theorem 3.1), paying a negligible price in exchange for the ability to re-use the calibration data. Second, compared to a full-conformal-style prediction band that applies the partition mechanism on the union of the calibration set and the test data points (and thereby would have to be re-fitted for all values of the training conformity score), our robust tree approach allows fitting the partition model only once. Our computational approach is different from a related approach by Gibbs et al. [18] who perform conditional quantile estimation over certain function classes, avoiding refitting on augmented data by exploiting monotonicity properties and dual formulation of quantile regression.
@@ 2 Conformal Tree
The core idea behind our self-grouping conformal prediction method is to learn the group information by applying a tree model to the conformity scores of the calibration dataset (Section 2.1). Since this ordinarily may violate the exchangeability assumption between the test and calibration data points, as such a tree would only be fit on the calibration data but not on the test point, we develop a tree algorithm that is robust to adding a new test observation (Section 2.2).
@@ 2.1 Calibration by Self-grouping
Note that the prediction interval construction relies on the model’s performance in the overall input space. For example, with the absolute residual score $S(x,y)=|y-\hat{\mu}(x)|$, the set $\widehat{C}_{n}(X_{n+1})$ is equivalent to $[\hat{\mu}(X_{n+1})-\hat{q},\hat{\mu}(X_{n+1})+\hat{q}]$, where a constant $\hat{q}=S^{*}(\mathcal{D}_{1:n})$ is added and subtracted over the data domain $\mathcal{X}$.
To locally sharpen the prediction interval, we can consider adding and subtracting a constant that differs depending on the location of $X_{n+1}$. Assume that the data space $\mathcal{X}=\cup_{k=1}^{K}\mathscr{X}_{k}$ is partitioned into non-overlapping groups $\mathfrak{X}=\{\mathscr{X}_{1},\ldots,\mathscr{X}_{K}\}$. We allow for categorical covariates which naturally serve as a group indicator. If $X_{n+1}\in\mathscr{X}_{k}$, we can calibrate the conformal interval based on only data points inside $\mathscr{X}_{k}$ to obtain a group-conditional coverage guarantee. In other words, defining
@@
Our method, called Conformal Tree, is described inside Algorithm 1. This procedure has an additional partition step compared to group-conformal inference [18, 3] because we do not assume that the grouping information is known a priori. We apply a regression tree on the conformity scores of the calibration dataset $\mathcal{D}_{1:n}$ using a potentially high-dimensional vector of both continuous and discrete covariates. Denote by $\widehat{\mathcal{T}}(\mathcal{D}_{1:n})\in\mathbb{T}_{K_{\rm max}}$ the tree fitted on $\mathcal{D}_{1:n}$, where $\mathbb{T}_{K_{\rm max}}$ is the class of binary dyadic trees with maximum ${K_{\rm max}}$ leaf nodes (defined formally later). The partition induced by the tree $\widehat{\mathcal{T}}(\mathcal{D}_{1:n})$ is denoted by $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n})$, where calibration is performed locally through (2.1) with $\mathfrak{X}=\widehat{\mathfrak{X}}(\mathcal{D}_{1:n})$. The fact that grouping information depends only on the calibration data, and not the test point, can potentially lead to a violation of group-conditional exchangeability for the conformity scores (see, Lemma 2). However, if we were to obtain a partition $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$ based on the union of calibration and testing data $\mathcal{D}_{1:n+1}$, the triples $\left(X_{1},Y_{1},S(X_{1},Y_{1})\right),...,\left(X_{n+1},Y_{n+1},S(X_{n+1},Y_{n+1})\right)$ would still be exchangeable when conditioned on $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$. Motivated by this observation, we propose a grouping mechanism based on a robust tree approach (Section 2.2), which guarantees that the partition does not change too often after adding one more observation, i.e. $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n})=\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$ with high probability. Such "unchangeability" property guarantees that the conditional coverage (that would be guaranteed if $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$ were used for splitting) is inherited with high probability when $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n})$ is used instead.
@@
1:Fitted model $\hat{\mu}$, calibration dataset $\mathcal{D}_{1:n}$, tree hyperparameters ($m,{K_{\rm max}},\eta=0.05$)
@@
Remark 1.
We highlight that Conformal Tree also identifies relevant subsets of the covariate space in which conditional coverage is guaranteed. The tree-based nature of group creation endorses identifiable and explainable groups which can often be understood in terms of some physical meaning. For example, one could identify that a model has more inherent uncertainty (and thus larger prediction sets) when classifying members of a specific race and gender combination. Refer to Figure 7 where this is demonstrated on a skin disease classification example.
@@ 2.2 Robust Regression Trees
We focus on dyadic trees where splits along covariates may only be placed at specific predetermined locations (dyadic rationals for continuous predictors rescaled to a unit interval). In general, our implementation allows for discrete predictors, provided that users define a way of splitting them. For example, we can map discrete covariates to a subset of the real line and split this component of the domain as if they were continuous.
We define a binary tree $\mathcal{T}$ as a collection of hierarchically organized nodes $(l,k)$ such that $(l,k)\in\mathcal{T}\Rightarrow(j,\lceil k/2^{l-j}\rceil)\in\mathcal{T}$ for $j=0,...,l-1$. Here $l$ denotes the depth, $k$ denotes the horizontal position, and the root node is defined to be $(0,0)$. The left and right children of a node $(l,k)$ are given by $(l+1,2k)$ and $(l+1,2k+1)$, respectively. A node $(l,k)\in\mathcal{T}$ is called an external or leaf node if it has no children. We denote the set of external nodes of $\mathcal{T}$ by $\mathcal{T}_{\rm leaves}$ and the internal nodes $\mathcal{T}_{\rm int}=\mathcal{T}\setminus\mathcal{T}_{\rm leaves}$. The internal nodes are assigned a split direction while the leaf nodes are not. We denote these split directions by $\mathcal{H}(\mathcal{T})=\{h_{lk}\in\{1,\dots,d\}:(l,k)\in\mathcal{T}_{int}\}$. We consider only dyadic trees, for which the tree nodes correspond to axis-aligned dyadic splits. We define $\mathbb{T}_{{K_{\rm max}}}$ as a space of all dyadic binary trees that split at most ${K_{\rm max}}-1$ times in one of the $d$ directions yielding up to ${K_{\rm max}}$ leaves. The partition $\mathfrak{X}(\mathcal{T})$, induced by both $\mathcal{T}$ and $\mathcal{H}(\mathcal{T})$, consists of regions associated with the leaf nodes $\mathcal{T}_{\rm leaves}$. In particular, each leaf node $(l,k)\in\mathcal{T}_{\rm leaves}$ corresponds to a rectangular region of the input space $\mathcal{X}$, which we denote as $\mathcal{X}_{(l,k)}(\mathcal{T},\mathcal{H}(\mathcal{T}))$. Given data $X_{1},\ldots,X_{n}\in\mathcal{X}$ and scores $S_{1},\ldots,S_{n}\in\mathbb{R}$, we denote the subset of data points that fall inside the $(l,k)^{th}$ region as $X_{(l,k)}(\mathcal{T},\mathcal{H}(\mathcal{T}))=\{X_{i}:X_{i}\in\mathcal{X}_{(l,k)}(\mathcal{T},\mathcal{H}(\mathcal{T}))\}$ and, similarly, $S_{(l,k)}(\mathcal{T},\mathcal{H}(\mathcal{T}))=\{S_{i}:X_{i}\in\mathcal{X}_{(l,k)}(\mathcal{T},\mathcal{H}(\mathcal{T}))\}$.
Our robust tree-fitting criterion is based on a splitting rule that targets the response range (as opposed to deviance). At every split opportunity, we split at the terminal node in the direction that yields the largest range decrease. Below, we formally define the range reduction of a node $(l,k)$ if it were split dyadically along the dimension $j$ denoted by $I_{R}^{j}(l,k)$.
@@
For a tree $\mathcal{T}\in\mathbb{T}_{K_{\rm max}}$ with split directions $\mathcal{H}(\mathcal{T})$, the range of a node $(l,k)\in\mathcal{T}_{\rm leaves}$ is given by $R_{lk}(\mathcal{T})=\max S_{(l,k)}(\mathcal{T},\mathcal{H}(\mathcal{T}))-\min S_{(l,k)}(\mathcal{T},\mathcal{H}(\mathcal{T}))\,.$ When a node $(l,k)$ splits along the $j^{\textrm{th}}$ direction, we define the range of the left child as $R^{j,\rm left}_{lk}(\mathcal{T})=\max S_{(l+1,2k)}(\widetilde{\mathcal{T}},\mathcal{H}^{j})-\min S_{(l+1,2k)}(\widetilde{\mathcal{T}},\mathcal{H}^{j})\,,$ where $\widetilde{\mathcal{T}}=\mathcal{T}\cup\{(l+1,2k),(l+1,2k+1)\}$ and $\mathcal{H}^{j}=\mathcal{H}(\mathcal{T})\cup\{h_{lk}=j\}.$ The range of the right child $R^{j,\rm right}_{lk}(\mathcal{T})$ is defined similarly as $S_{(l+1,2k+1)}(\widetilde{\mathcal{T}},\mathcal{H}^{j})$. Then, the range reduction obtained by splitting a node $(l,k)$ in dimension $j$ is given by
$I_{lk}^{j}(\mathcal{T})=R_{lk}(\mathcal{T})-\left[R^{j,\rm left}_{lk}(\mathcal{T})+R^{j,\rm right}_{lk}(\mathcal{T})\right]\,.$ (2.2)
As is typical with CART implementations [40], we regularize the tree and define a stopping criterion by the following three hyperparameters: the minimal number of samples per leaf $m$, maximum number of leaves ${K_{\rm max}}$, and minimum range reduction rate $\eta=0.05$. A node is called a candidate node if there exists a direction along which a split would not cause the tree to violate any of these criteria.
@@
We use the notation from Definition 2.1. For a tree $\mathcal{T}\in\mathbb{T}_{K_{\rm max}}$ such that $|\mathcal{T}_{\rm leaves}|<{K_{\rm max}}$, the set of eligible split directions for a node $(l,k)\in\mathcal{T}_{\rm leaves}$ is defined as
@@
We continue to sequentially add splits to the tree by maximizing the range reduction among candidate nodes until either we reach the maximum number of leaves ${K_{\rm max}}$, or there are no candidates remaining, at which time we stop the tree fitting procedure.
@@
1:Dataset $\mathcal{D}=\{(X_{i},S_{i}):i\in[n]\}$, minimum samples per leaf $m$, max leaves ${K_{\rm max}}$
@@
3:while $|\mathcal{T}_{\mathrm{leaves}}|<{K_{\rm max}}$  do
@@
The new robust tree algorithm can be extended to incorporate the $X_{n+1}$ value (not $y_{n+1}$) for tree fitting. While this variation gives tighter theoretical bounds, it introduces non-negligible additional computational cost that scales linearly with the number of test points. For this reason, we defer it to Section B in the Supplement.
@@
Our robust tree algorithm resembles greedy CART [4]. Similarly, we require a minimum improvement (range reduction rate) and a minimum number of observations per leaf. Rather than using cross-validation to prune back, we rely on early stopping once we reach ${K_{\rm max}}$ leaf nodes or until splitting would yield less than $m$ observations per leaf node, or an insufficient improvement. The tuning parameter ${K_{\rm max}}$ is pre-determined, unlike Bayesian implementations of CART [10] which are both locally and globally adaptive. We use a fixed value of 0.05 for the minimum improvement threshold for all of the experiments in this paper.